A car with bad shock absorbers bounces up and down after hitting a bump. The car has a mass of 2500 kg and is supported by four springs, each having a spring constant of 6000 N/m. What is the period for each spring? 


1
Expert's answer
2021-10-13T12:02:30-0400

"k_{total}=k+k+k+k=4k=4\cdot6000=24000\ (N/m)"


"T=2\pi\sqrt{m/k}=2\cdot 3.14\cdot\sqrt{2500/24000}=2.03\ (s)" . Answer